Dehesa, Inc. has 8,000 shares of 5%, $15 par, cumulative preferred stock and 50,000 shares of $3 par common stock outstanding. No dividends were declared last year. However, the board of directors just declared a $34,000 dividend this year. What amount of the total dividend will be paid to common stockholders?

$34,000

$22,000

$28,000

$12,000
Answer #1

Preferred dividends for last year = (8,000*15*5%) = 6,000

Preferred dividends for current year = (8,000*15*5%) = 6,000

Dividends paid to common stockholders = 34,000 - (6,000+6,000)

= 22,000